OpenStudy (anonymous):

Use the MacLaurin series for f(x). Compute the 10th derivative of f(x)=arctan(x^2/4) at x=0. f^(10)(0)=

OpenStudy (amistre64):

well, the 0 will cancel out all the x parts; and your left with the constant

OpenStudy (amistre64):

\[\frac{d^n}{dx^n}x^a=aPn\ x^{a-n}\]if i see it correctly

OpenStudy (anonymous):

arctan(x)=x-x^3/3+x^5/5-...+\[(-1)^{k}x ^{2k+1}/(2k+1)\]

OpenStudy (amistre64):

\[y = tan^{-1}(x^2/4)\] \[tan(y)=x^2/4\] \[sec^2(y)=x/2\] \[y'=x cos^2(y)/2\] \[cos^2(y)=\frac{16}{x^2+16}\] \[y' = \frac{8x}{x^2+16}\] gotta dbl chk that to be sure tho

OpenStudy (amistre64):

x^4 not x^2 :)

OpenStudy (amistre64):

16+x^4 = 16(\(1+\frac{x^4}{16}\)) 8x/2^0-x^5/2^1+x^9/2^5 - x^13/2^9 ------------- \(1+\frac{x^4}{16}\) ) 8x (8x+x^5/2) ----------- -x^5/2 (-x^5/2-x^9/2^5) ------------- x^9/32 (x^9/32+x^13/2^9) \[16\left(8x+\sum_{inf}^{n=1}\frac{(-1)x^{4n+1}}{2^{4n-3}}\right) \] maybe ....

OpenStudy (amistre64):

i spose 8x/16 = x/2 tho so id have to adjust for that

OpenStudy (anonymous):

that's so complicated;(

OpenStudy (amistre64):

lol, im just trying to brush up on somethings with this really

OpenStudy (anonymous):

i knew an example like this. \[e ^{-x ^{2}}=\sum_{k=0}^{\infty} (-1)^{k}x ^{2k}/(k!)\] =1-x^2+x^4/2!-x^6/3!+x^8/4!+... \[f ^{7}(0)=0\] \[f ^{8}(0)/8!=1/4!\] \[f ^{8}(0)=8!/4!\]

OpenStudy (anonymous):

but when i applied to my question,it is not correct,not sure whether it is a calculation problem or...

OpenStudy (anonymous):

\[arctanx=\sum_{0}^{\infty}(-1)^{k}x ^{2k+1}/(2k+1)\]

OpenStudy (amistre64):

im trying to get the power series that is the derivative of arctan(x^2/4) so that I can determine the power series of arctan(x^2/4) thru integration of the polynomial

OpenStudy (amistre64):

\[\frac{d}{dx}tan(x^2/4)=\frac {8x}{16+x^4}\to \ \frac{x}{2+\frac{x^4}{2^3}}\] \[\frac{x}{2+\frac{x^4}{2^3}}=\sum_{n=0}^{inf }\frac{(-1)^nx^{4n+1}}{4n+1}\] \[\int \sum_{n=0}^{inf }\frac{(-1)^nx^{4n+1}}{4n+1} dx = \sum_{n=0}^{inf }\frac{(-1)^nx^{4n+2}}{(4n+1)(4n+2 )}=tan^{-1}(\frac{x^2}{4})\]

OpenStudy (anonymous):

is it you are right?

OpenStudy (amistre64):

well, i got the power series for the 1st derivative right :)

OpenStudy (anonymous):

i think the answer should be 40320. x^2/2-x^6/30+x^10/90 f^(10)(0)/10!=1/90

OpenStudy (amistre64):

forgot to type in the denominators 2^(4n+1)

OpenStudy (amistre64):

im sure it could very well be; but i havent gotten that far yet :)

OpenStudy (anonymous):

but in this case,it will be equal to 0!!

OpenStudy (amistre64):

\[f'(x)=g(x)=\sum_{n=0}^{inf }\frac{(-1)^nx^{4n+1}}{4n+1}\] \[\frac{d}{dx}x^k=k x^{k-1}\] \[\frac{d^2}{dx^2}x^k=k(k-1) x^{k-2}\] \[\frac{d^3}{dx^3}x^k=k(k-1)(k-2) x^{k-3}\] \[\frac{d^n}{dx^n}x^k=^kP_n\ x^{k-n}\] \[\frac{d^n}{dx^n}\left( g(x)=\sum_{n=0}^{inf }\frac{(-1)^n}{4n+1}x^{4n+1} \right)\]

OpenStudy (amistre64):

the 9th derivative of g(x) is the 10th derivative of f(x) :)

OpenStudy (amistre64):

all that typing and the site crashes on me :/

OpenStudy (amistre64):

\[\frac{d^9}{dx^9}\ g(x)=c_0\ ^{1}P_{9}\ x^{1-9}-c_1\ ^{5}P_{9}\ x^{5-9}+c_2\ ^{9}P_{9}\ x^{9-9}...\] \[\frac{d^9}{dx^9}\ g(x)= c_2\ ^{9}P_{9}+0...=9.8.7.6.5.4.3.2.1\frac{1}{2^9}\]

OpenStudy (amistre64):

9.8.7.6.5.4.3.2.1 222 22 2 9.7.6.5.3.1 ------------ 2.2.2 9.7.3.5.3 --------- 2.2 81*35/4 maybe :)

OpenStudy (amistre64):

woohoo!! http://www.wolframalpha.com/input/?i=10th+derivative+arctan%28x%5E2%2F4%29 2835/4 = 81*35/4

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

:)